A rectangular parking lot has a length that is 6 yards greater than the width. The area of the parking lot is 720 square yards. Find the length and width.
Answer by Cromlix(4381)

Width = x
Length = x + 6
Area = Length * Width
720 yds^2 = x(x + 6)
x^2 + 6x = 720
x^2 + 6x - 720 = 0
(x + 30)(x - 24) = 0
x = -30 (No answer)
x = 24
Width = 24yds
Length = 30yds
Area = 30 * 24 = 720 yds^2
Hope this helps.
:-)